OpenStudy (anonymous):

The midpoint of CD is E(-1, 0) One endpoint is C(5.2). What are the coordinates of the other endpoint?

OpenStudy (kirbykirby):

The midpoint coordinates are: \[ \left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2} \right)=(-1, 0)\] You are given one point, say \((x_1, y_1)=(5, 2)\) So for the x-coordinate of the midpoint, you can write the following: \[ \frac{x_1+x_2}{2}=\frac{5+x_2}{2}=-1\] For the y-coordinate midpoint, you have \[ \frac{y_1+y_2}{2}=\frac{2+y_2}{2}=0\]

OpenStudy (kirbykirby):

can you solve for \(x_2\) and \(y_2\) in each case?

OpenStudy (anonymous):

yes, thank you(:

OpenStudy (kirbykirby):

\(\Huge \checkmark\) :)